# Diagnosing and Addressing Pitfalls in KG-RAG Datasets: Toward More Reliable Benchmarking

### v2

@@ 1 Introduction
To better understand the limitations of existing benchmarks, we conduct a detailed manual inspection of 16 publicly available KGQA datasets including the widely adopted WebQSP and CWQ yih2016value ; talmor2018web ; bao2016constraint ; su2016generating ; ngomo20189th ; zhang2018variational ; azmy2018farewell ; saha2018complex ; trivedi2017lc ; dubey2019lc ; jiang2019freebaseqa ; keysers2019measuring ; gu2021beyond ; perevalov2022qald ; cao2020kqa ; dammu2025dynamic . A detailed summary of these datasets, along with our sampling and evaluation protocol, is provided in Appendix B. Our analysis reveals several recurring issues that compromise their utility for evaluating KG-RAG systems. These include factually incorrect or outdated ground truth answers, and ambiguously phrased or trivially simple questions. In particular, WebQSP and CWQ, which serve as the dominant [^2] Around 30 papers on KG-RAG released between 2023 and 2025 adopted WebQSP and/or CWQ for evaluation, highlighting their widespread use. A detailed breakdown is provided in Appendix B evaluation benchmarks in recent KG-RAG research xu2024generate ; gao2025frag ; xu2025memory ; wang2025reasoning ; zhang2025trustuqa , exhibit serious deficiencies. Specifically, we found only 52% of sampled WebQSP examples and 49.3% of CWQ examples to be factually correct. Across all 16 datasets, the average correctness rate is just 57%, based on manual evaluation of over 1,000 question–answer pairs. A detailed summary of dataset quality is provided in Table 1. Additionally, many datasets rely on rigid exact-match metrics that penalize semantically correct answers expressed in alternative forms, further limiting their reliability risch2021semantic ; pan2021attacking ; wang2023evaluating .
@@ 3.1 Dataset Quality Issues
To assess the reliability of existing KGQA benchmarks, we manually inspected over 1,000 question-answer pairs sampled from 16 widely used datasets. A summary of findings is provided in Table 1. A description of the datasets and inspection protocol is included in appendix B. For each dataset, we randomly selected a representative subset of examples and evaluated them along three key dimensions: (1) factual correctness of the annotated answer, and (2) clarity and appropriateness of the question. We paid particular attention to the WebQSP and CWQ datasets, as they are the most dominant benchmarks in recent KGQA research. As shown in Table 1, for WebQSP and CWQ, we sampled 100 and 300 examples, respectively. For the remaining datasets, we uniformly sampled 60 examples. Overall, our inspection revealed substantial quality issues across many benchmarks. Notably, the most widely used WebQSP and CWQ datasets show serious quality issues. In WebQSP, only 52% of the sampled answers were judged correct, with quality issues including inaccurate answer and poor questions. Similarly, CWQ, which builds on WebQSP as its seed, achieved a correctness rate of just 49.33%, suffering from similar flaws along with additional issues due to increased question complexity. Beyond these two, we found that over half of the evaluated datasets had correctness rates below 60%, suggesting widespread challenges with annotation accuracy and question clarity. Next, we discuss the specific issues.
@@ 4 KGQAGen: A Framework for Grounded KGQA Dataset Construction
The overall process of KGQAGen is illustrated in Figure 1. The framework consists of three main stages: (1) Seed Subgraph Initialization: The process begins by selecting a seed entity and constructing a local subgraph by retrieving related facts from the KG. This subgraph provides the initial context for reasoning. (2) Question Generation through Iterative LLM-Guided Subgraph Expansion:
To support non-trivial, multi-hop questions, we iteratively expand the subgraph by traversing neighboring entities and relations. This involves alternating between KG traversal and LLM evaluation. At each step, the subgraph is expanded by exploring neighboring entities and relations within the KG. After each expansion, an LLM is prompted to evaluate whether the subgraph contains sufficient information to support a well-formed, multi-hop question. If not, the expansion continues. Once the subgraph is judged sufficient, the LLM generates a natural language question, identifies the corresponding answer set, extracts a minimal supporting subgraph, and constructs the associated SPARQL query. (3) Answer Validation and Refinement: The final stage ensures that the generated answer set is correct and fully grounded in the KG. To achieve this, the generated SPARQL query is executed against the knowledge base to retrieve the actual answers. If the results match the generated answer set, the instance is accepted.
Before we proceed to detail the framework, we introduce the notation used throughout this section.
@@ 5 KGQAGen-10k: A Sample Dataset Generated by KGQAGen
Manual Quality Assessment. To assess the factual accuracy of KGQAGen-10k, we conducted a manual audit of 300 randomly sampled question–answer pairs. Following the same annotation protocol used in Section 3, each example was carefully verified. We found that 289 out of 300 examples (96.3% ) are correct. This result suggests that the QA instances generated by KGQAGen are highly reliable and well-grounded in verifiable knowledge. We defer analysis of question difficulty to Section 6, where we benchmark several LLM and KG-RAG models. A case study of the generated questions is provided in Appendix E.
Statistics and Properties. We analyze KGQAGen-10k along two key dimensions: linguistic complexity and topic coverage: (1) Linguistic Complexity: Most questions (61.1%) are 16–30 words long, showing moderate complexity. A third exceed 30 words, indicating deeper reasoning. Only 7.5% are short factoid queries, highlighting the dataset’s focus on rich, natural questions. For answer set size, 84.5% of the questions yield a single answer, while 9.7% return three or more answers. (2) Topic Coverage: The dataset also provides broad semantic coverage. Figure 2 shows that KGQAGen-10k covers a broad range of topics. The most represented categories are Arts (42.3%) and Astronomy (17.3%), followed by STEM fields (16% combined), Sports, Geography, and Philosophy.
@@ 6 KGQAGen-10k as a Benchmark for KG-RAG Models
Experiment Setup. We benchmark all models on KGQAGen-10k using a standardized split of 8,629 training, 1,079 development, and 1,079 test examples. Most prior work evaluates KGQA systems using an exact string match between predicted and reference answers. However, as highlighted in Section 3, this metric often fails to capture semantically valid predictions that differ in surface form. To address this limitation, we introduce LLM-Assisted Semantic Match (LASM), a lightweight verification mechanism that activates only when a prediction fails the exact match. LASM uses GPT-4o-mini to assess semantic equivalence between predictions and gold answers (details in Appendix F). We report Accuracy, Hit@1, Precision, Recall, and F1 under both Exact Match (EM) and LASM for direct comparison.
Evaluated Models. We evaluate three categories of models on KGQAGen-10k: (1) Pure LLMs: This group includes a range of open-source and commercial models—LLaMA-3.1-8B-Instruct, LLaMA2-7B, Mistral-7B-Instruct-v0.2, GPT-4o-mini, GPT-4, GPT-4o, GPT-4.1, and Deepseek-Chat. These models receive only the natural language question without any retrieval or external grounding; (2) KG-RAG Models: We include recent KG-RAG models such as RoG, GCR, ToG , and PoG. These models use the KG as an external retrieval source, either by incorporating retrieved triples, augmenting the prompt with symbolic paths, or integrating topology-aware context; and (3) LLM with Supporting Subgraph (LLM-SP). To assess performance under perfect retrieval, we include LLaMA2-7B and GPT-4o models that are directly given the associated supporting subgraph of the questions from the dataset construction process. These setups simulate perfect retrieval and help isolate the reasoning capabilities from retrieval effectiveness. See Appendix F for model details.
@@ B.2 Recent KG-RAG Publications with WebQSP and/or CWQ
WebQSP [70] and CWQ [56] have emerged as the primary benchmarks for empirical evaluation in LLM-based knowledge graph question answering. Table 3 presents a chronological summary of recent LLM-based KGQA models, indicating for each whether WebQSP and CWQ were used for evaluation and providing a brief description of the model’s main approach. Notably, nearly 30 KG-ARG models released between 2023 and 2025 have adopted one or both datasets as central components of their experimental protocols, despite the quality issues identified in Section 3.
@@ C.1.1 Inaccurate Ground Truth Answers
  • ID: CWQ-848_ac67410188d0f2258139a3c84773885e
    Question: What is the time zone in the location where the time zone is Central Western?
    Answer: Parliamentary system, Constitutional monarchy, Federal monarchy
    Issue: The answers are not time zones but forms of government. This is an incorrect annotation, and the question is self-answering and non-informative.
@@ C.1.2 Low-Quality or Ambiguous Questions
  • ID: CWQ-80_3dafd01d90b628c5913e0c64c1143354
    Question: Who inspired the famous person who went to Willem II College?
    Answer: Eugène Delacroix, Claude Monet, Jean-François Millet, Jozef Israëls, etc.
    Issue: The query is open-ended; it does not specify which famous person.
@@
  • ID: CWQ-848_ac67410188d0f2258139a3c84773885e
    Question: What is the time zone in the location where the time zone is Central Western?
    Answer: Parliamentary system, Constitutional monarchy, Federal monarchy
    Issue: The question is self-answering, non-informative, and offers little reasoning challenge.
@@
  • ID: CWQ-2621_7360e892294860c6ef7ad9a10e540e1b
    Question: Which author wrote editions of "Notes from My Travels ’s husband"?
    Answer: Brad Pitt
    Issue: The question is ill-formed and refers to a non-existent book.
  • ID: CWQ-1304_f0c99f377c8b944364e85b70b9f9b331
    Question: Which education institution is Bundesrealgymnasium Linz the leader of?
    Answer: Hitler Youth, Gestapo, Nazi Party, etc.
    Issue: Ill-formed query; Bundesrealgymnasium Linz is itself an educational institution, not a leader of other entities.
@@ D.2 SPARQL Validation Prompt
Appendix E A Case Study of KGQAGen-10k
@@
Appendix F Experimental Details
@@ F.1 Model Specifications
F.2 Beyond Exact Match: Introducing LASM

### v4

@@ 1 Introduction
To better understand the limitations of existing benchmarks, we conduct a detailed manual inspection of 16 publicly available KGQA datasets including the widely adopted WebQSP and CWQ yih2016value ; talmor2018web ; bao2016constraint ; su2016generating ; ngomo20189th ; zhang2018variational ; azmy2018farewell ; saha2018complex ; trivedi2017lc ; dubey2019lc ; jiang2019freebaseqa ; keysers2019measuring ; gu2021beyond ; perevalov2022qald ; cao2020kqa ; dammu2025dynamic . A detailed summary of these datasets, along with our sampling and evaluation protocol, is provided in Appendix B. Our analysis reveals several recurring issues that compromise their utility for evaluating KG-RAG systems. These include factually incorrect or outdated ground truth answers, and ambiguously phrased or trivially simple questions. In particular, WebQSP and CWQ, which serve as the dominant [^2] Around 30 papers on KG-RAG released between 2022 and 2025 adopted WebQSP and/or CWQ for evaluation, highlighting their widespread use. A detailed breakdown is provided in Appendix B evaluation benchmarks in recent KG-RAG research xu2024generate ; gao2025frag ; xu2025memory ; wang2025reasoning ; zhang2025trustuqa , exhibit serious deficiencies. Specifically, we found only 52% of sampled WebQSP examples and 49.3% of CWQ examples to be factually correct. Across all 16 datasets, the average correctness rate is just 57%, based on manual evaluation of over 1,000 question–answer pairs. A detailed summary of dataset quality is provided in Table 1. Additionally, many datasets rely on rigid exact-match metrics that penalize semantically correct answers expressed in alternative forms, further limiting their reliability risch2021semantic ; pan2021attacking ; wang2023evaluating .
@@ 3.1 Dataset Quality Issues
To assess the reliability of existing KGQA benchmarks, we manually inspected over 1,000 question-answer pairs sampled from 16 widely used datasets. A summary of findings is provided in Table 1. A description of the datasets and inspection protocol is included in appendix B. For each dataset, we randomly selected a representative subset of examples and evaluated them along three key dimensions: (1) factual correctness of the annotated answer, and (2) clarity and appropriateness of the question and (3) current evaludation of exact match shortcoming. We paid particular attention to the WebQSP and CWQ datasets, as they are the most dominant benchmarks in recent KGQA research. As shown in Table 1, for WebQSP and CWQ, we sampled 100 and 300 examples, respectively. For the remaining datasets, we uniformly sampled 60 examples. Overall, our inspection revealed substantial quality issues across many benchmarks. Notably, the most widely used WebQSP and CWQ datasets show serious quality issues. In WebQSP, only 52% of the sampled answers were judged correct, with quality issues including inaccurate answer and poor questions. Similarly, CWQ, which builds on WebQSP as its seed, achieved a correctness rate of just 49.33%, suffering from similar flaws along with additional issues due to increased question complexity. Beyond these two, we found that over half of the evaluated datasets had correctness rates below 60%, suggesting widespread challenges with annotation accuracy and question clarity. Next, we discuss the specific issues.
@@ 4 KGQAGen: A Framework for Grounded KGQA Dataset Construction
The overall process of KGQAGen is illustrated in Figure 1. The framework consists of three main stages: (1) Seed Subgraph Initialization: The process begins by selecting a seed entity and constructing a local subgraph by retrieving related facts from the KG. This subgraph provides the initial context for reasoning. (2) Question Generation through Iterative LLM-Guided Subgraph Expansion: To support non-trivial, multi-hop questions, we iteratively expand the subgraph by traversing neighboring entities and relations. This involves alternating between KG traversal and LLM evaluation. At each step, the subgraph is expanded by exploring neighboring entities and relations within the KG. After each expansion, an LLM is prompted to evaluate whether the subgraph contains sufficient information to support a well-formed, multi-hop question. If not, the expansion continues. Once the subgraph is judged sufficient, the LLM generates a natural language question, identifies the corresponding answer set, extracts a minimal supporting subgraph, and constructs the associated SPARQL query. (3) Answer Validation and Refinement: The final stage ensures that the generated answer set is correct and fully grounded in the KG. To achieve this, the generated SPARQL query is executed against the knowledge base to retrieve the actual answers. If the results match the generated answer set, the instance is accepted. An ablation study presented in Appendix E further confirms the importance of each component in KGQAGen. Both LLM-guided subgraph expansion and SPARQL-based verification are shown to be essential for producing accurate and well-grounded QA pairs. Before we proceed to detail the framework, we introduce the notation used throughout this section.
@@ 5 KGQAGen-10k: A Sample Dataset Generated by KGQAGen
Manual Quality Assessment. To assess the factual accuracy of KGQAGen-10k, we conducted a manual audit of 300 randomly sampled question–answer pairs. Following the same annotation protocol used in Section 3, each example was carefully verified. We found that 289 out of 300 examples (96.3% ) are correct. This result suggests that the QA instances generated by KGQAGen are highly reliable and well-grounded in verifiable knowledge. We defer analysis of question difficulty to Section 6, where we benchmark several LLM and KG-RAG models. A case study of the generated questions is provided in Appendix F.2.
Statistics and Properties. We analyze KGQAGen-10k along two key dimensions: linguistic complexity and topic coverage: (1) Linguistic Complexity: Most questions (61.1%) are 16–30 words long, showing moderate complexity. A third exceed 30 words, indicating deeper reasoning. Only 7.5% are short factoid queries, highlighting the dataset’s focus on rich, natural questions. For answer set size, 84.5% of the questions yield a single answer, while 9.7% return three or more answers. (2) Structural Difficulty: Following DyVal zhu2023dyval , we compute six graph-based metrics for each supporting subgraph—number of nodes, edges, average degree, depth (hops), width (maximum frontier), and extra links (non-chain shortcuts). Across all 10,787 examples, 98% require 2–5 hops, 84% contain 5–30 entities, 83% have 4–28 relations, 79% exhibit width 4–20, and 92% are fully connected. These statistics confirm that KGQAGen-10k consists of diverse, non-trivial reasoning graphs with substantial structural complexity. A full summary table is provided in Appendix F.1. (3) Topic Coverage: The dataset also provides broad semantic coverage. Figure 2 shows that KGQAGen-10k covers a broad range of topics. The most represented categories are Arts (42.3%) and Astronomy (17.3%), followed by STEM fields (16% combined), Sports, Geography, and Philosophy.
@@ 6 KGQAGen-10k as a Benchmark for KG-RAG Models
Experiment Setup. We benchmark all models on KGQAGen-10k using a standardized split of 8,629 training, 1,079 development, and 1,079 test examples. Most prior work evaluates KGQA systems using an exact string match between predicted and reference answers. However, as highlighted in Section 3, this metric often fails to capture semantically valid predictions that differ in surface form. To address this limitation, we introduce LLM-Assisted Semantic Match (LASM), a lightweight verification mechanism that activates only when a prediction fails the exact match. LASM uses GPT-4o-mini to assess semantic equivalence between predictions and gold answers (details in Appendix G). We report Accuracy, Hit@1, Precision, Recall, and F1 under both Exact Match (EM) and LASM for direct comparison.
Evaluated Models. We evaluate three categories of models on KGQAGen-10k: (1) Pure LLMs: This group includes a range of open-source and commercial models—LLaMA-3.1-8B-Instruct, LLaMA2-7B, Mistral-7B-Instruct-v0.2, GPT-4o-mini, GPT-4, GPT-4o, GPT-4.1, and Deepseek-Chat. These models receive only the natural language question without any retrieval or external grounding; (2) KG-RAG Models: We include recent KG-RAG models such as RoG, GCR, ToG , and PoG. These models use the KG as an external retrieval source, either by incorporating retrieved triples, augmenting the prompt with symbolic paths, or integrating topology-aware context; and (3) LLM with Supporting Subgraph (LLM-SP). To assess performance under perfect retrieval, we include LLaMA2-7B and GPT-4o models that are directly given the associated supporting subgraph of the questions from the dataset construction process. These setups simulate perfect retrieval and help isolate the reasoning capabilities from retrieval effectiveness. See Appendix G for more details of baseline models.
@@
Cross-Dataset Comparison. To further contextualize the results, we evaluate representative KG-RAG models on two established KGQA benchmarks, WebQSP (yih2016value, ) and CWQ (talmor2018web, ), in addition to our KGQAGen-10k. While models such as RoG, GCR, ToG, and PoG achieve high Hit@1 scores on WebQSP and CWQ (e.g., 85.7 and 92.2, respectively), their performance drops sharply on KGQAGen-10k (21.3–54.0), reflecting its higher reasoning complexity and stricter grounding requirements. Detailed results are provided in Appendix G.3. This consistent decline across models highlights that legacy benchmarks often overestimate true reasoning ability, whereas KGQAGen-10k offers a more challenging and diagnostic testbed for evaluating KG-RAG systems.
@@ B.2 Recent KG-RAG Publications with WebQSP and/or CWQ
WebQSP [73] and CWQ [58] have emerged as the primary benchmarks for empirical evaluation in LLM-based knowledge graph question answering. Table 3 presents a chronological summary of recent LLM-based KGQA models, indicating for each whether WebQSP and CWQ were used for evaluation and providing a brief description of the model’s main approach. Notably, nearly 30 KG-RAG models released between 2022 and 2025 have adopted one or both datasets as central components of their experimental protocols, despite the quality issues identified in Section 3.
@@ C.1.1 Inaccurate Ground Truth Answers
  • ID: CWQ-1182_ac67410188d0f2258139a3c84773885e
    Question: Who is the current queen of the location whose time zone is Central Western Time Zone?
    Answer: Elizabeth II
    Issue: The answer is outdated. The Central Western Time Zone (UTC+8:45) corresponds to a small region in Australia. Since Queen Elizabeth II passed away in 2022, the current monarch of Australia is King Charles III.
@@
  • ID: CWQ-124_7360e892294860c6ef7ad9a10e540e1b
    Question: What movie did the author who published editions for Notes from My Travels direct?
    Answer: Unbroken
    Issue: The annotation is incomplete. The book Notes from My Travels was written by Angelina Jolie, who has directed multiple films, including In the Land of Blood and Honey (2011), Unbroken (2014), By the Sea (2015), and First They Killed My Father (2017).
@@ D.2 SPARQL Validation Prompt
Appendix E Ablation Study on KGQAGen Design
Setup. We perform an ablation study to quantify the contribution of core components in the KGQAGen-10k generation pipeline. In our framework, questions are generated through iterative LLM-guided subgraph expansion combined with symbolic SPARQL-based verification. To isolate the effect of each component, we consider three configurations: (A1) random subgraph selection replacing LLM-guided expansion, (A2) alternative generation LLMs, and (A3) inclusion or exclusion of SPARQL-based answer validation. A3 is treated as a cross-setting axis applied to all variants. We randomly sample 100 seed entities and report two metrics: generation success rate (percentage of seeds producing valid questions) and end-to-end success rate (percentage producing fully validated QA pairs).
Setting Generation Rate (%) End-to-End Rate (%) Validation Effect (%)
Full KGQAGen (GPT-4.1) 94 77 +17
A1: Random Subgraph 100 62 +15
A2-1: LLaMA-3-70B 57 35 +4
A2-2: GPT-4o-mini 65 41 +6
Table 4: Ablation results on 100 randomly selected seed entities. “Validation Effect” denotes the improvement in valid QA pairs after applying SPARQL verification.
Findings. The full KGQAGen pipeline achieves a 94% generation success rate and a 77% end-to-end validation rate. Removing LLM-guided subgraph selection (A1) maintains full generation coverage but lowers validation to 62%, as random sampling often introduces extraneous entities that cause ambiguity. To assess the effect of model choice, we replace GPT-4.1 with smaller LLMs under the same setup. LLaMA-3-70B generates 57 questions with 35 validated (61% of generated), while GPT-4o-mini produces 65 questions with 41 validated (63% of generated). Smaller models frequently simplify multi-hop reasoning to single-hop relations or omit relational constraints. In contrast, the full KGQAGen pipeline with GPT-4.1 yields 82 generated and 77 validated questions, achieving both higher coverage and correctness. These results demonstrate that model capacity and structure-aware subgraph guidance jointly determine data quality, while SPARQL-based symbolic verification mitigates hallucinations and constrains potential biases.
To better understand the failure modes of symbolic verification, we analyzed 420 instances filtered out by the SPARQL-based validation stage. The main error types include: empty query results (166 cases, 39.5%), excessively large answer sets (184 cases, 43.8%), and answer mismatches where the annotated answer was absent from the SPARQL results (70 cases, 16.7%). These issues primarily stem from incomplete or ambiguous question–answer pairs, limited KG coverage, or challenges in SPARQL query formulation. Future work will strengthen the validation module with tighter constraints, clarification prompts, and finer-grained error classification to improve explainability and systematically address validation failures.
Appendix F KGQAGen-10k Analysis
F.1 KGQAGen-10k Statistics
Structural Complexity Metrics. Following DyVal [81], we measure six structural aspects of the supporting subgraphs for all 10,787 examples in KGQAGen-10k: (1) number of nodes, (2) number of edges, (3) average degree, (4) reasoning depth (hops), (5) width (maximum frontier size), and (6) extra links (non-chain connections). These metrics collectively capture both the scale and topological difficulty of multi-hop reasoning paths.
Measure Distribution Range Percentage of Examples
Depth (hops) 2–5 98%
Nodes (entities) 5–30 84%
Edges (relations) 4–28 83%
Width (max frontier) 4–20 79%
Reachability Fully connected 92%
Table 5: Graph-based reasoning complexity analysis for KGQAGen-10k following DyVal-style metrics. The dataset exhibits diverse and non-trivial structural properties, demonstrating its suitability for evaluating reasoning over complex knowledge graphs.
Across these dimensions, most questions involve 2–5-hop reasoning chains over compact yet densely connected subgraphs. The high connectivity (92%) and wide range of entity and relation counts illustrate that KGQAGen-10k contains structurally rich and diverse reasoning contexts, providing a challenging evaluation setting for KG-RAG and KGQA models.
F.2 A Case Study of KGQAGen-10k
@@
Appendix G Experimental Details
@@ G.1 Model Specifications
G.2 Beyond Exact Match: Introducing LASM
@@
To ensure evaluation reliability, LASM is applied only when the prediction fails the exact match, serving as a selective fallback rather than a replacement for EM. This design mitigates potential risks of LLM-based scoring such as hallucinations or factual drift, as LASM is invoked for only a small subset of cases. To assess its accuracy, we manually inspected 100 randomly sampled instances where LASM was triggered. Two human reviewers independently verified semantic correctness using Wikipedia, identifying only one false positive (1% error rate)—a case where the prediction “Austronesian languages” was judged equivalent to the ground truth “Oceanic”. This low error rate demonstrates that LASM provides a reliable supplement to exact match evaluation. Furthermore, similar approaches in the NLG domain support the robustness of LLM-based evaluation, as shown by recent studies such as G-Eval [31], FActScore [37], MT-Bench [80], and QAFactEval [68], which report strong alignment between GPT-4-based judgments and human evaluation.